How far apart are two conducting plates that have an electric field strength of 5.30×104 V/m between them, if their potential difference is 18.0 kV?
Answer #1

We know that the formula is V=E*d where V=potential difference, E=electric field, d= spacing

now d=V/E =18000/5.30*10^4= 0.3396 m= 33.96 cm